OpenStudy (anonymous):

divide (4x^3-26x+25) / (2x-4)

OpenStudy (jiteshmeghwal9):

is this 4x^2-26x+25 ?

OpenStudy (anonymous):

no 4x^3

OpenStudy (anonymous):

the rest is right

OpenStudy (anonymous):

(4x^3-26x+25) / (2x-4) = 2x^2 2x fits in 4x^3 , 2x^2 times. which makes 4*2x^2 = 6x^2 so then it's: (-6x^2-26x+25) / (2x-4) = 2x^2 Then continue from here.

OpenStudy (anonymous):

PS 2x^2 is not the answer, just the start of the answer.

OpenStudy (anonymous):

would it be \[2x ^{2}+4x-5+\frac{ 5 }{ 2x-4 }\]

OpenStudy (anonymous):

yes, except i got 5/(2x-2) Just a sec. Have to doublecheck

OpenStudy (anonymous):

Still got 5/(2x-2) on the last bit rest is correct.